# Language Modeling for Fast Inference

### v1

@@ Abstract
This paper presents the Block Transformer architecture which adopts hierarchical global-to-local modeling to autoregressive transformers to mitigate the inference bottlenecks of self-attention. To apply self-attention, the key-value (KV) cache of all previous sequences must be retrieved from memory at every decoding step. Thereby, this KV cache IO becomes a significant bottleneck in batch inference. We notice that these costs stem from applying self-attention on the global context, therefore we isolate the expensive bottlenecks of global modeling to lower layers and apply fast local modeling in upper layers. To mitigate the remaining costs in the lower layers, we aggregate input tokens into fixed size blocks and then apply self-attention at this coarse level. Context information is aggregated into a single embedding to enable upper layers to decode the next block of tokens, without global attention. Free of global attention bottlenecks, the upper layers can fully utilize the compute hardware to maximize inference throughput. By leveraging global and local modules, the Block Transformer architecture demonstrates 10–20x gains in inference throughput compared to vanilla transformers with equivalent perplexity. Our work introduces a new approach to optimize language model inference through novel application of global-to-local modeling.
@@ 1 Introduction
Generating tokens with transformer-based autoregressive language models (LMs) is costly due to the self-attention mechanism that attends to all previous tokens [6, 66]. To alleviate the cost of the self-attention, it is common to cache the key-value (KV) states of all tokens across all layers during the autoregressive decoding. However, while each decoding step only computes the KV state of a single token, it still has to load the KV states of all previous tokens for computing self-attention scores. Subsequently, this KV cache IO mostly dominates the inference cost spent on serving LMs. While several techniques have been proposed for reducing the inference cost of the attention component [20, 35, 69], developing effective transformer-based LM architectures that inherently avoid the attention overhead is still an ongoing challenge.
Hierarchical global-to-local architectures [49, 31] have shown significant potential to effectively model large-scale data by addressing global dependencies in coarse detail and capturing fine details within local regions. Inspired by these frameworks, we identify a unique opportunity to mitigate key bottlenecks in autoregressive transformer inference: (1) coarse global modeling can reduce overall costs by its granularity; but more importantly, (2) localized self-attention can nearly eliminate the costs of attention as there is no need to compute, store, and retrieve KV-cache of past tokens beyond the small local context.
This paper presents the Block Transformer architecture which models global dependencies through self-attention between coarse blocks (each representing multiple tokens) at lower layers, and decodes fine-grained tokens within each local block at upper layers, as shown in Figure 1. Specifically, a lightweight module called (1) the embedder first embeds each block of $L_{B}$ input tokens into an input block embedding. These become the input units of (2) the block decoder, an autoregressive transformer that applies self-attention between blocks to decode a context block embedding which contains information for predicting the next block. Finally, (3) the token decoder autoregressively decodes the token contents of the next block, applying local self-attention between only the $L_{B}$ tokens within the block. While this leaves the token decoder to solely rely on the output block embedding for global context information, it drastically reduces self-attention costs to be linear to total context length, and eliminates the need to prefill prompt tokens during inference.
While analogous transformer architectures have been proposed to handle long sequences comprised of raw bytes [74], prior work consider the global module to be the primary model, benefiting from coarse processing, while the embedder and local module simply map between coarse and fine representations to reduce context length. Our approach to global-to-local modeling in LMs challenges these prior beliefs, and uncovers substantial inference-time benefits that have been overlooked in previous work. In detail, we propose that both the global block decoder and local token decoder can play vital roles in language modeling, hence the term global-to-local language modeling. Our ablations reveal that a more balanced parameter allocation across the global and local modules enhances performance, and also results in higher throughput due to significantly shortened context lengths in the local module.
Figure 1: An overview of the Block Transformer architecture, demonstrated with a block length of four (each alphabet symbol represents one token from the vocabulary). The shaded parts indicate prompt tokens, which do not need to be prefilled for the token decoder during inference. The receptive field of the last token is illustrated with a green line, demonstrating how global-to-local language modeling efficiently covers the full context in the receptive field.
@@
Extensive experiments on models up to 1.4 billion parameters show that Block Transformers notably improve inference throughput for both prefill- and decode-intensive scenarios, achieving 10–20$\times$ gains in throughput compared to vanilla transformers with equivalent perplexity or zero-shot task performance. Despite the architectural restriction of global attention in Block Transformers, our models show similar ability to utilize global context compared to their vanilla transformer counterparts. In addition, we show that it is possible to uptrain pretrained vanilla model into Block Transformers, closely approaching the performance of those pretrained from scratch, using just 10% of the training budget for adaptation.
Our main contributions are summarized below:
  • We are the first to recognize the central role and inference-time benefits of both global and local modeling in autoregressive transformers–particularly the significance of local modules.
  • We leverage these insights to optimize inference throughput in our architecture to significantly extend the Pareto frontier of performance to throughput compared to vanilla transformers.
@@ 2 Block Transformer
The Block Transformer employs global and local attention mechanisms with hierarchical paradigm by separating the comprehension of the full context and detailed interactions into two distinct stages. Precisely, global context is captured at lower layers as coarse block-level granularity, where each block consists of a fixed number of tokens aggregated into a single embedding. The local dependencies are resolved at upper layers, where multiple subword tokens are decoded in an autoregressive manner by solely attending context block embedding from the block decoder. The Block Transformer consists of three components:
@@
2.1 Why is Block Transformer efficient?
The main goal of our architecture design is to minimize the wall-clock bottlenecks during inference. In vanilla transformers, the global treatment of self-attention to all previous tokens significantly hinders batch decoding throughput, mainly due to memory overhead of retrieving previous KV cache [20, 25]. This also necessitates all prompt tokens, which are typically quite lengthy, to be fully prefilled prior to decoding the first token, contributing to increased latency [1, 25].
A global-to-local approach can mitigate these costs by isolating the expensive bottlenecks of global modeling to the lower layers and perform local modeling within independent blocks at the upper layers. Coarse-grained global modeling (block-level decoding) alleviates KV cache bottlenecks by a factor of block length, while maintaining the ability to account for the full context Local decoding comes free of the cost of prefill, and nearly removes KV cache overhead, thus benefits from significantly higher utilization of the compute units on inference hardware. This allows the token decoder to use more FLOPs for fine-grained language modeling with minimal impact on inference throughput. Table 1 outlines the principal wall-time bottlenecks at the prefill and decode stages, and summarizes the efficiency gains of our block and token decoders.
Although our models require more parameters than vanilla transformers to maintain comparable performance, the actual bottleneck in throughput is the KV cache overhead, allowing our model to still achieve higher speed improvements. Thereby, we focus on production systems like cloud platforms, which can accommodate the higher parameter demands. Edge devices are constrained by memory [3] and typically use small batches [61]. Since parameter IO is a critical bottleneck [51], we leave the optimization of the Block Transformer for on-device scenarios to future work.
Table 1: Comparison of relative compute and memory costs for our block and token decoder compared to vanilla transformers, and overview of principal bottlenecks for each inference stage. The number of layers is represented by $N$, the dimension by $D$, the batch size by $B$, the context length by $L$, and the block length by $L_{B}$. $\dagger$ Token decoder is not used during the prefill stage, so its complexity is zero. The details about inference efficiency are summarized in Appendix E.
Inference bottleneck Relative complexity
Metric Name Prefill Decode Vanilla Block $\text{Token}^{\dagger}$
Memory Parameter - - $\mathcal{O}(ND^{2})$ $1$ $1$
KV cache - - $\mathcal{O}(NDBL)$ $1/L_{B}$ $L_{B}/L\approx 0$
IO Parameter $\mathcal{O}(ND^{2}L)$ $1/L_{B}$ $1$
KV cache   $\mathcal{O}(NDBL^{2})$ $1/L_{B}^{2}$ $L_{B}/L\approx 0$
FLOPs Attention score   $\mathcal{O}(NDBL^{2})$ $1/L_{B}^{2}$ $L_{B}/L\approx 0$
Linear projections   $\mathcal{O}(ND^{2}BL)$ $1/L_{B}$ $1$
@@ 2.2 Embedder
Our embedder design prioritizes simplicity given the small block length (2–8) in our study. We primarily use a lookup table $E_{\text{emb}}\in\mathbb{R}^{V\times D_{\text{emb}}}$ to retrieve and concatenate trainable token embeddings, where the token embedding dimension $D_{\text{emb}}$ is set to $D/L_{B}$, with $D$ being the dimension of block representations used throughout the network). While we explored variants such as small encoder transformers (Appendix F), these did not yield performance improvements (subsection 3.4).
@@ 2.3 Block decoder
The block decoder aims to contextualize block representations by attending to preceding blocks, utilizing the embedder’s output as input. This autoregressive transformer operates at the block level, producing output block embeddings (also called context embeddings) that enable the token decoder to autoregressively decode the subsequent block’s token contents. Given input block embeddings from the embedder, derived from input tokens $x_{0:(i\times L_{B}-1)}$, the block decoder outputs a context embedding which contains the information to predict $x_{(i\times L_{B}):((i+1)\times L_{B}-1)}$.
This approach mitigates the quadratic costs of self-attention by using coarse-grained block inputs instead of individual tokens, while preserving global modeling capabilities and ease of hardware acceleration of dense attention [75]. This reduces the context length of a given sequence by $L_{B}$ compared to a vanilla transformer. In terms of FLOPs (the main bottleneck during prefill), all positionwise computations are reduced by a factor of $L_{B}$, and attention score computation is reduced by $L_{B}^{2}$ [74]. During decoding, KV cache usage and KV cache IO (the main bottleneck during batch decoding) are reduced by $L_{B}$ and $L_{B}^{2}$, respectively, allowing for larger batch sizes and higher compute utilization.
@@ 2.4 Token decoder
The token decoder locally decodes the individual tokens of the next block using the context block embedding as the sole source of global context information. The token decoder is also a standard autoregressive transformer, featuring it’s own embedding table $E_{\text{tok}}\in\mathbb{R}^{V\times D_{\text{tok}}}$ and classifier. The key to designing the token decoder lies in how to incorporate the context embedding into the decoding process, in a way that effectively leverages the high compute density of the token decoder.
The token decoder eliminates prefill (necessary only in the block decoder), as context information is provided by the output block embedding–hence the term context embedding. Additionally, KV cache IO, a major bottleneck during batch decoding, is nearly removed. While vanilla attention’s KV cache IO is quadratic to the full context length ($L^{2}$), the token decoder’s local attention costs $L_{B}^{2}$ per block over $L/L_{B}$ blocks, resulting in a linear cost to the full context length and a reduction factor of $L/L_{B}$ (e.g., $2048/4=256$ in our main models). This allows for significantly higher compute unit utilization compared to vanilla transformers, which have ~1% model FLOPs utilization (MFU) [51], making the inference wall-time cost of extra FLOPs relatively cheap.
To incorporate the context embedding and leverage this low-cost compute, we project the context block embedding into prefix tokens, enabling further refinement of the global context. Expanding the number of prefix tokens (prefix length) broadens the token decoder’s computation width and allows for finer attention to context information, similar to pause tokens [29]. Owing to parallel processing and small local context, these extra prefix tokens do not incur significant wall-time overhead. While we also considered summation and cross-attention based variants (Appendix F), these proved less effective than our main method (subsection 3.4).
@@ 3.2 Main results
The actual inference throughput and memory efficiency of the Block Transformer are significantly higher compared to vanilla models. We measure the maximum throughput [60], which use maximum batch sizes of each model variant allowed by memory. As shown in 2(a) and 2(b), our models achieve Pareto-optimality, especially demonstrating up to 25 times increase, under two scenarios: prefill-heavy and decode-heavy, where the input and output sequence lengths are 2048, 128 and vice-versa. This efficiency improvement is due to effective reductions in KV cache memory, which allows batch sizes to be about six times larger, as summarized in memory per sample in Table 2. The Block Transformer further reduces latency in a prefill-heavy setting, as past KV states of prompts need to be cached only in the block decoder, without forwarding them to the token decoder.
The Pareto frontiers for variable fixed batch sizes, i.e., 1, 32, and 256, are illustrated in Appendix I. We discover that as both the model size and batch size increase, the throughput rate of the Block Transformer scales exponentially. Considering that the LLMs typically utilized in real-world applications have billions of parameters, and taking into account the strategy of aggregating multiple user requests to optimize batch inference [35, 50, 60], the results suggest that our proposed architecture will demonstrate even more benefits in practical multi-tenant deployment scenarios.
@@ 3.3 Analysis on parameter allocation ratio and block length
We measure average loss at each position within a block, depicted in 3(d). The position-wise loss typically exhibits a U-shaped pattern, aligning with findings from a previous multiscale language model [74] and blockwise parallel decoding methods [62, 14, 34]. This trend stems from the lack of global context in context embeddings, which escalates uncertainty at later positions. Moreover, perplexity at specific positions correlates with the parameter sizes of two decoders. A larger block decoder significantly lowers initial position loss due to predictions solely based on the context embedding. In contrast, a larger token decoder improves prediction accuracy for later tokens by better leveraging local context. These interdependent effects dictate the optimal parameter ratio, with similar patterns evident in models of various sizes, detailed in Appendix J.
@@
3(b) demonstrates that training loss still follows a U-shaped pattern across different allocation ratios, regardless of block length. Optimal ratios shift with block length: shorter blocks benefit from a larger block decoder, while longer blocks perform better with more parameters in the token decoder. This is due to the inverse relationship between block length and FLOPs of the block decoder, which influences model capacity [22, 23, 29]. As 3(e) shows, first position loss significantly decreases with shorter blocks, reflecting increased capacity in the block decoder. While the token decoder shows minimal differences in FLOPs across block lengths, it has more chance to improve the likelihood of later tokens as block length increases, favoring a larger token decoder. These trends are consistent across different model sizes and allocation ratios, detailed in Appendix K.
@@
We evaluate the allocation ratio and block length from a throughput perspective, summarizing the Pareto frontier in Appendix L. Models with larger token decoders reach Pareto-optimality by achieving higher throughput at a minor performance compromise. Since KV cache IO significantly influences inference time, allocating more parameters to the token decoder is advantageous because the local context length is bounded by the block length. Additionally, increasing the block length improves throughput as KV cache length in the block decoder reduces proportionally. Therefore, although our main configuration uses a one-to-one ratio and a block length of four, opting for a longer block length and a larger token decoder could result in a higher-throughput model.

### v2

@@ Abstract
We introduce the Block Transformer which adopts hierarchical global-to-local modeling to autoregressive transformers to mitigate the inference bottlenecks associated with self-attention. Self-attention requires the key-value (KV) cache of all previous sequences to be retrieved from memory at every decoding step to retrieve context information, leading to two primary bottlenecks during batch inference. First, there is a significant delay in obtaining the first token, as the information of the entire prompt must first be processed to prefill the KV cache. Second, computation of subsequent tokens is bottlenecked by the high memory I/O demand of fetching the entire KV cache, which grows linearly with sequence length, incurring quadratic memory reads overall. We design the Block Transformer to strategically mitigate these costs, by incorporating coarsity and locality into an integrated global-to-local architecture. At the lower layers, we aggregate tokens into fixed size blocks to apply attention across the entire sequence at coarse-grained detail, to capture the global context while minimizing KV cache overhead. At upper layers, we apply attention within each block to decode individual tokens, to model fine-grained details with a lightweight local KV cache. We pretrain vanilla and Block Transformers from scratch and demonstrate that Block Transformers reach 10–20x inference throughput compared to vanilla transformers with equivalent perplexity and zero-shot task performance.
@@ 1 Introduction
Generating tokens with transformer-based autoregressive language models (LMs) is costly due to the self-attention mechanism that attends to all preceding tokens [6, 77]. To minimize computation, it is common to cache the key-value (KV) states of all tokens during decoding. However, significant initial latency remains from processing the KV states of all prompt tokens during the first decoding step. Also, while subsequent decoding steps only need to compute a single token, this is often bottlenecked by the memory access required to retrieve the KV states of all previous tokens. While numerous techniques have been proposed to reduce attention costs [23, 42, 2], there has been limited research on effective architectures that structurally mitigate attention overheads in transformer-based LMs.
Hierarchical architectures [59, 36, 22, 56, 54, 57, 28, 50, 84] have shown potential in efficiently modeling long sequences such as character-level text or pixel-level images by pooling inputs into coarser units. While most of these employ upsampling to regress back to a finer level for attention computation, several approaches [50, 84] further enhance efficiency through local processing within each pooled unit. Nevertheless, they have not recognized or explored the potential of applying local processing to alleviate the overheads of autoregressive inference, explained above.
In this paper, we present the Block Transformer architecture which employs a global-to-local hierarchical structure at the subword-level to holistically maximize the inference benefits of global-to-local processing for self-attention. We begin with a brief summary of our architecture as shown in Figure 1:
The lower layers are designed to efficiently capture the global input context by applying attention at the coarse level. First, the lightweight embedder maps each block of $L_{B}$ subword tokens into a single input block embedding. These are passed to the block decoder, an autoregressive transformer that applies self-attention between blocks to output a output block embedding, or context embedding, used by the upper layers to decode the next block. The upper layers are designed to predict the next block in finer detail, returning to the level of individual tokens. Here, the token decoder autoregressively decodes the tokens in the next block. Relying on the context embedding from the lower layers for global context, fine-level attention is applied only within the current block of $L_{B}$ tokens, using a minuscule local KV cache.
Figure 1: An overview of the Block Transformer architecture with a block length of $L_{B}=4$. Each letter represents a standard subword token. Attention is restricted within the dotted boundaries in the decoders. The KV values outside the boundary of the currently decoded block do not need to be retrieved during decoding, and can be discarded from memory. Consequently, given the prompt tokens, A, B, …, L, the shaded parts do not need to be computed in the token decoder.
This structure alleviates a wide range of key inference bottlenecks with only simple modifications to the standard transformer. At the lower layers, coarse processing reduces the amount of KV cache storage by a factor of $L_{B}$ and memory access by ${L_{B}}^{2}$. At the upper layers, attention is applied within the local block of $L_{B}$ tokens, nearly eliminating its overhead. Consider a prompt with $L$ tokens. While vanilla transformer layers would have to prefill and store the KV cache of $L$ tokens and retrieve these at every decoding step, the upper layers of the Block Transformer only need to see up to $L_{B}\ll L$ most recent tokens, nearly eliminating prefill computation and KV cache overheads.
While prior work [84] has introduced similar structures, it has largely overlooked their potential benefits in autoregressive inference. Notably, they focus on optimizing overall FLOPs for efficient pretraining by exploiting lightweight local modules that simply map between coarse and fine representations. Our approach challenges this viewpoint, uncovering vital roles of both the global block decoder and local token decoder in language modeling. Ablations reveal that a more balanced parameter allocation across the global and local modules enhances performance and inference throughput, attributed to significant reduction in self-attention overheads within the local module.
@@
Extensive experiments on models up to 1.4 billion parameters show that Block Transformers notably improve inference throughput for both prefill- and decode-intensive scenarios, achieving 10–20$\times$ gains in throughput compared to vanilla transformers with equivalent perplexity or zero-shot task performance. Despite the architectural restriction of global-to-local attention, our models show a comparable ability to utilize full context on recent long-context benchmarks, such as PG19 [62] and Needle-In-a-Haystack [39], when compared to their vanilla transformer counterparts. In addition, we show that it is possible to uptrain pretrained vanilla models into Block Transformers, closely approaching the performance of those pretrained from scratch, using minimal training budget.
@@ 2 Block Transformer
The Block Transformer employs global and local attention mechanisms with hierarchical paradigm by separating the comprehension of the full context and detailed interactions into two distinct stages. Specifically, global context is captured at lower layers in coarse block-level granularity, where each block consists of a fixed number of tokens aggregated into a single embedding. The local dependencies are resolved at upper layers, where multiple subword tokens are decoded in an autoregressive manner using the context embedding from the block decoder, with local attention.
The Block Transformer consists of three components:
@@
We outline the design and list efficiency benefits for each component in the following subsections. To lay the groundwork for detailed cost analysis, we begin with a simplified primer on key bottlenecks in autoregressive transformers, assuming a single accelerator device.
2.1 A primer on the key bottlenecks of autoregressive transformers
Key costs of autoregressive transformers
These can be categorized into compute, parameter memory and KV cache memory. (1) Compute refers to arithmetic operations, dominated by matrix multiplications in the attention and feedforward layers. The number of floating-point operations (FLOPs) is proportional to the number of non-embedding parameters and total tokens, i.e., sequence length $\times$ batch size. (2) Parameters must be fully stored in device memory and retrieved at each forward or backward pass, regardless of the sequence length and batch size. (3) During inference, the KV cache of previously computed sequences are typically cached in memory, and retrieved at each forward pass, i.e., decoding step. Similar to compute, its size is proportional to sequence length $\times$ batch size.
Case study of Llama 7B [75]
The compute cost of each token is roughly 7 $\times$ 2 = 14GFLOPs, where 2 comes from the multiply-accumulate operation used in matrix-vector multiplication [40]. The memory size of the model parameters is 7 $\times$ 2 = 14GB under 16-bit, or $2$-byte, floating-point precision. The size of the KV cache of a single token is 512KB. While this seems minuscule compared to the parameters, the total KV cache can quickly outsize the parameters with more tokens; for sequence length 2048 and batch size 16, the KV cache occupies 16GB. To translate this to wall-clock time, note that the compute throughput (FLOP/s) of accelerator devices is 2–3 orders of magnitude higher than HBM memory bandwidth (bytes/s)—a gap which is widening exponentially [33].
Computation scenarios for autoregressive LMs
These can be broadly divided into training and inference. During training, every token is processed in parallel. For a precise cost analysis of inference, we further dissect it into the prefill stage and the decode stage. To generate a response to a given prompt, all input tokens must first be processed to obtain and cache their KV values in each layer, allowing subsequent tokens to access them for self-attention. This is referred to as the prefill stage. Next is the decode stage, where subsequent tokens are generated one at a time. In each forward step, only one token per batch sample is computed, but the KV cache of all preceding tokens must be loaded from memory.
Primary bottlenecks per computation scenario vary significantly
(1) During training, all tokens are processed in parallel, thus compute demands far outweigh parameter memory access, which remains constant. (2) Similarly, the prefill stage is typically compute-bound, as all input tokens are processed in parallel. (3) In contrast, the decode stage, given sufficient sequence length, is heavily memory-bound, as only one token per batch sample is computed in each forward pass, while parameter and KV cache memory access demands are significant. Specifically, KV cache memory access exceeds parameter memory access when sequence length and batch size are large, while parameter memory dominates when they are small. A larger batch size helps reduce the relative cost of parameter memory access by amortizing it over batch samples. Our proposed Block Transformer design optimizes inference, especially in batch decoding, where KV cache impacts throughput.
@@ 2.2 Embedder
For the embedder, we prioritize simplicity, given the small block lengths $L_{B}={1,2,4,8}$ considered in our study. Our primary design uses a lookup table $E_{\text{emb}}\in\mathbb{R}^{V\times D_{\text{emb}}}$ to retrieve and concatenate trainable token embeddings. We set the embedding dimension to $D_{\text{emb}}=D/L_{B}$, where $D$ is the primary model dimension, used throughout the block and token decoders. We also consider variants which incorporate small encoder transformers (Appendix F), but these do not yield performance improvements (subsection 3.4).
@@ 2.3 Block decoder
The block decoder aims to contextualize block representations by attending to preceding blocks, utilizing the embedder’s output as input. This autoregressive transformer operates at the block level, producing output block embeddings, or context embeddings [^1] We use the term context embedding, as it is the sole source of context information for the token decoder., that enable the token decoder to autoregressively decode the subsequent block’s token contents. Given input block embeddings from the embedder, derived from input tokens $x_{0:i\times L_{B}-1}$, the block decoder outputs a context embedding which contains the information to predict $x_{i\times L_{B}:(i+1)\times L_{B}-1}$.
This approach mitigates the quadratic costs of self-attention by using coarse-grained block inputs instead of individual tokens, while preserving global modeling capabilities and ease of hardware acceleration of dense attention [85]. This reduces the context length of a given sequence by $L_{B}$ compared to a vanilla transformer. We compare the block decoder with vanilla transformer layers, in terms of the key bottlenecks discussed in subsection 2.1:
  • Computation is reduced by $L_{B}$ due to reduced input units.
  • Parameter memory access is reduced by $L_{B}$ due to reduced decoding frequency.
  • KV cache size is reduced by $L_{B}$ due to coarsity. [^2] Reduced KV cache memory footprint can lower hardware requirements or allow for larger batch sizes, which can increase throughput by further amortizing parameter memory access. KV cache access is reduced by $L_{B}^{2}$ due to reduced size and decoding frequency.
@@ 2.4 Token decoder
The token decoder decodes the individual tokens of the next block, taking the form of an autoregressive transformer with a separate embedding table $E_{\text{tok}}\in\mathbb{R}^{V\times D_{\text{tok}}}$ and classifier. Each block is processed independently, relying on the context embedding as the sole source of information on previous input blocks. Only the local KV cache of the current block needs to be stored in memory and accessed at each decoding step. Conveniently, none of the prompt tokens in previous blocks are attended by future tokens, allowing the prefill stage to be skipped entirely–except for the most recent block.
[…]
  • Computation is reduced to near-zero during prefill. Training computes remains the same.
  • Parameter memory access remains the same.
  • KV cache size is reduced by $R=L/L_{B}$ with $L_{B}\ll L$, practically eliminating its memory footprint.3 KV cache memory access is equally reduced by $R$.
The key to designing the token decoder lies in how to incorporate the context embedding into the decoding process. In our main design, we project the context embedding into prefix tokens, enabling further refinement of the global context. Expanding the number of prefix tokens, i.e., prefix length, broadens the token decoder’s computation width and allows for finer attention to context information and improved performance, similar to pause tokens [34]. Note that this is only feasible due to local processing, whereas extra tokens in vanilla transformers layers would exacerbate the overhead of KV cache. Extra computation incurred by prefix tokens has minimal effect on inference throughput as inference is largely memory-bound. While we also considered summation and cross-attention based variants (Appendix F), these proved less effective than our main method (subsection 3.4).
@@ 3.2 Main results
The actual inference throughput and memory efficiency of the Block Transformer are significantly higher compared to vanilla models. We measure the maximum throughput [71], which use maximum batch sizes of each model variant allowed by memory. As shown in 2(a) and 2(b), our models achieve Pareto-optimality, especially demonstrating up to 25 times increase, under two scenarios: prefill-heavy and decode-heavy, where the input and output sequence lengths are 2048, 128 and vice-versa. This efficiency improvement is due to effective reductions in KV cache memory, which allows batch sizes to be about six times larger, as summarized in memory per sample in Table 1. The Block Transformer further reduces latency in a prefill-heavy setting, as past KV states of prompts need to be cached only in the block decoder, without forwarding them to the token decoder. As detailed in Appendix I, this overall trend remains consistent even with the FlashAttention algorithm [23] employed during inference.
The Pareto frontiers for variable fixed batch sizes, i.e., 1, 32, and 256, are illustrated in Appendix J. We discover that as both the model size and batch size increase, the throughput rate of the Block Transformer scales exponentially. Considering that the LLMs typically utilized in real-world applications have billions of parameters, and taking into account the strategy of aggregating multiple user requests to optimize batch inference [42, 60, 71], the results suggest that our proposed architecture will demonstrate even more benefits in practical multi-tenant deployment scenarios.
@@ 3.3 Analysis on parameter allocation ratio and block length
We measure average loss at each position within a block, depicted in 3(d). The position-wise loss typically exhibits a U-shaped pattern, aligning with findings from a previous multiscale language model [84] and blockwise parallel decoding methods [72, 16, 41]. This trend stems from the lack of global context in context embeddings, which escalates uncertainty at later positions. Moreover, perplexity at specific positions correlates with the parameter sizes of two decoders. A larger block decoder significantly lowers initial position loss due to predictions solely based on the context embedding. In contrast, a larger token decoder improves prediction accuracy for later tokens by better leveraging local context. These interdependent effects dictate the optimal parameter ratio, with similar patterns evident in models of various sizes, detailed in Appendix K.
@@
3(b) demonstrates that training loss still follows a U-shaped pattern across different allocation ratios, regardless of block length. Optimal ratios shift with block length: shorter blocks benefit from a larger block decoder, while longer blocks perform better with more parameters in the token decoder. This is due to the inverse relationship between block length and FLOPs of the block decoder, which influences model capacity [25, 26, 34]. As 3(e) shows, first position loss significantly decreases with shorter blocks, reflecting increased capacity in the block decoder. While the token decoder shows minimal differences in FLOPs across block lengths, it has more chance to improve the likelihood of later tokens as block length increases, favoring a larger token decoder. These trends are consistent across different model sizes and allocation ratios, detailed in Appendix L.
@@
We evaluate the allocation ratio and block length from a throughput perspective, summarizing the Pareto frontier in Appendix M. Models with larger token decoders reach Pareto-optimality by achieving higher throughput at a minor performance compromise. Since KV cache IO significantly influences inference time, allocating more parameters to the token decoder is advantageous because the local context length is bounded by the block length. Additionally, increasing the block length improves throughput as KV cache length in the block decoder reduces proportionally. Therefore, although our main configuration uses a one-to-one ratio and a block length of four, opting for a longer block length and a larger token decoder could result in a higher-throughput model.